Two boxes each contain red and white colored bulbs. The percentage of red bulbs in the first box is 30% and 10% in the second box. One box is chosen at random and a bulb is withdrawn.

a. Construct a probability tree and show all relevant probabilities.
b. What is the probability of withdrawing a red bulb?
c. If a red bulb is withdrawn, what is the probability that it came from box 2?
1
Expert's answer
2019-10-10T11:03:22-0400

Let "I" be the event that the first box is chosen, "II" be the event that the second box is chosen, "R" be the event that the red bulb is chosen, and "W" be the event that the white bulb is chosen.

Given that


"P(I)=P(II)=0.5, P(R\ | \ I)=0.3, P(R\ |\ II)=0.1"

Construct a probability tree and show all relevant probabilities. 

We have that


"P(W\ | \ I)=1-P(R\ | \ I)=0.7,""P(W\ | \ II)=1-P(R\ |\ II)=0.9"

Then


"P(I\cap R)=P(I)P(R\ | \ I)=0.5(0.3)=0.15""P(I\cap W)=P(I)P(W\ | \ I)=0.5(0.7)=0.35""P(II\cap R)=P(II)P(R\ | \ II)=0.5(0.1)=0.05""P(II\cap W)=P(II)P(W\ | \ II)=0.5(0.9)=0.45"


b. What is the probability of withdrawing a red bulb?

By the Theorem of total probability


"P(R)=P(I)P(R\ | I)+P(II)P(R\ | \ II)"

"P(R)=0.5(0.3)+0.5(0.1)=0.2"

c. If a red bulb is withdrawn, what is the probability that it came from box 2?

By the Bayes' rule


"P(II \ | \ R)={P(II)P(R\ | \ II) \over P(I)P(R\ | I)+P(II)P(R\ | \ II)}"

"P(II \ | \ R)={0.5(0.1) \over 0.5(0.3)+0.5(0.1)}=0.25"